OpenStudy (anonymous):

You are going to purchase a new car, but being a responsible consumer means doing a little bit of research first. First, you find the vehicle you are purchasing and its price. Vehicle: Chevy Volt Price: $39,145 Current interest rate: 3% Using the function A(t)=P(1+r/n)^nt, create the function that represents your new car loan that is compounded monthly. The principle will be the price of the vehicle you selected, not how much you are putting down. Help @SolomonZelman @midhun.madhu1987

OpenStudy (anonymous):

You are going to purchase a new car, but being a responsible consumer means doing a little bit of research first. First, you find the vehicle you are purchasing and its price. Vehicle: Chevy Volt Price: $39,145 Current interest rate: 3% 1. Using the function A(t)=P(1+r n )^nt, create the function that represents your new car loan that is compounded monthly. The principle will be the price of the vehicle you selected, not how much you are putting down. 2. Being a smart financial planner, you want to figure out how many months it will be until your principal is paid down to $10,000.00. Solve for t and show all of your work. Note that t will be negative because the number of months will decrease the principal. 3.Lastly, you decide to keep track of your loan four times a month instead of monthly. Solve for the adjusted interest rate. Remember to use the formula A(t)=P[(1+r n )1c cnt where c = 4. When solving for the adjusted interest rate, be sure to set it equal to 1+rn . Here's the problem

OpenStudy (anonymous):

Hey Beautiful you don't mind if you can send me some help? @sky577

OpenStudy (anonymous):

Hey @owl-lover @SolomonZelman @calculusxy @cwrw238 can you HELP PLEASE????

OpenStudy (anonymous):

@owl-lover @SolomonZelman @calculusxy @cwrw238 MEDAL WILL BE GIVEN

OpenStudy (anonymous):

the equation is: 39,145(1+0.03/12)^12t=10,000 t=-45.5 adjusted int. rate: (1+0.03/12)^3 =1.0075 rounded off =1.01

OpenStudy (anonymous):

@SCARLET_SPIDER

OpenStudy (anonymous):

Beautiful can you explain to me how you got t @sky577

OpenStudy (anonymous):

Take your time

OpenStudy (anonymous):

ok just wait a minute need to find the work i did on it

OpenStudy (anonymous):

Alright no problem Beautiful

OpenStudy (anonymous):

Your fine no problem